OpenStudy (anonymous):

If sin (A) = 5/13 with A in QII, find sin (A/2).

OpenStudy (anonymous):

Firstly you need to find cos(A) also..

OpenStudy (anonymous):

\[\cos(A) = \sqrt{1 - \sin^2(A)}\]

OpenStudy (anonymous):

Can you find \(\cos(A)\) ??

OpenStudy (anonymous):

sorry im here now

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{ -\sqrt{26} }{ 26}\]

OpenStudy (anonymous):

i think, im stuck here

OpenStudy (anonymous):

How you got \(\sqrt{26}\) there in numerator?

OpenStudy (anonymous):

multiplied by sqrt? is that wrong?

OpenStudy (anonymous):

Let us solve this part only first: \[1 - \frac{12}{13}\]

OpenStudy (anonymous):

Can you tell after solving what will you get?

OpenStudy (anonymous):

1/13

OpenStudy (anonymous):

Oh, I see now, where are you going.. Your answer is right in my opinion..

OpenStudy (anonymous):

but it was wrong

OpenStudy (anonymous):

Yeah it is wrong.

OpenStudy (anonymous):

your values of cos(A) is wrong I think.

OpenStudy (anonymous):

\[\cos(A) = \sqrt{1 - \sin^2(A)} = \sqrt{1 - \frac{25}{169}} = \sqrt{\frac{144}{169}} = \pm \frac{12}{13}\]

OpenStudy (anonymous):

Up to here, you got?

OpenStudy (anonymous):

I want to ask why have you chosen +12/13 for cos(A), As A is in second quadrant, and cos in second quadrant is negative, so don't you think you should use Negative 12/13 there?

OpenStudy (anonymous):

mistake

OpenStudy (anonymous):

Can you correct it now?

OpenStudy (anonymous):

-5\[-5\sqrt{26}/26?\]

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

what do i do?

OpenStudy (anonymous):

Choose: \[\cos(A) = -\frac{12}{13}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[1 + \frac{12}{13} = ??\]

OpenStudy (anonymous):

yes i did this already...

OpenStudy (anonymous):

\[\sin(\frac{A}{2}) = \sqrt{\frac{25}{26}}\]

OpenStudy (anonymous):

Now, tell me will you take positive value or negative value?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

\[\sin(\frac{A}{2}) = \sqrt{\frac{25}{26}} = \pm \sqrt{\frac{25}{26}}\]

OpenStudy (anonymous):

Okay, I agree with you, but why?

OpenStudy (anonymous):

sin is y and y is positive in qII